# Semilinear spaces of matrices with transitivity, rank or spectral constraints

### v1

@@ 5.6 The converse statement
Let $U$ and $V$ be finite-dimensional $\mathbb{D}$-vector spaces, with $\dim_{\mathbb{D}}U=\dim_{\mathbb{D}}V>0$, and set $n:=\dim_{\mathbb{D}}U$. Let $\mathcal{S}$ be an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$ such that $\dim_{\mathbb{F}}(\mathcal{S}x)=nd-1$ for all $x\in U\smallsetminus\{0\}$. Assume that $\mathcal{S}$ is target-reduced. Then all the nonzero alternators of $\mathcal{S}$ are nondegenerate, and $\operatorname{\operatorname{Alt}}(\mathcal{S})$ has dimension at most $1$.
@@
Let $b\in\operatorname{\operatorname{Alt}}(\mathcal{S})$ be nonzero and degenerate. It follows that $nd=\dim_{\mathbb{F}}V>1$. Every element of $\mathcal{S}$ maps the left radical $L$ of $b$ into its right radical $R$, and here $\dim_{\mathbb{F}}L=\dim_{\mathbb{F}}R$ since $\dim_{\mathbb{F}}U=\dim_{\mathbb{F}}V$.
We choose $x\in L\smallsetminus\{0\}$ and deduce that $\dim_{\mathbb{F}}R\geq\dim_{\mathbb{F}}\mathcal{S}x=nd-1$. Hence $\operatorname{\operatorname{rk}}(b)=1$. Then each bilinear form $(x,y)\in U^{2}\mapsto b(x,u(y))$, with $u\in\mathcal{S}$, is alternating with rank at most $1$, and hence is zero. Hence $\operatorname{\operatorname{Im}}u\subseteq R$ for all $u\in\mathcal{S}$, which contradicts the assumption that $\mathcal{S}$ is target-reduced. Hence, every nonzero element of $\operatorname{\operatorname{Alt}}(\mathcal{S})$ is nondegenerate.
Next, let $b$ and $b^{\prime}$ belong to $\operatorname{\operatorname{Alt}}(\mathcal{S})\smallsetminus\{0\}$. Choose $x\in U\smallsetminus\{0\}$. By the first step, $b(x,-)\neq 0$. The nonzero linear forms $b^{\prime}(x,-)$ and b(x,)b(x,-) vanish on the $\mathbb{F}$-linear hyperplane $\mathcal{S}x$, and hence $b^{\prime}(x,-)=\lambda b(x,-)$ for some $\lambda\in\mathbb{F}$. Then $b^{\prime}-\lambda b$ is a degenerate element in $\operatorname{\operatorname{Alt}}(\mathcal{S})$, and by the first step $b^{\prime}-\lambda b=0$. We conclude that $\dim\operatorname{\operatorname{Alt}}(\mathcal{S})\leq 1$. ∎
@@
Set $\mathcal{S}:=\mathcal{A}_{b,\mathbb{D}}$. To start with, to prove the first point we reduce the situation to the one where $B$ is nondegenerate. To see this, we consider the left radical $L$ of $B$, which equals the one of $b$, and the induced form $\overline{B}:(U/L)\times V\rightarrow\mathbb{D}$, and the $\mathbb{F}$-bilinear form $\overline{b}:(x,y)\in(U/L)\times V\mapsto e(\overline{B}(x,y))$. All the operators in $\mathcal{S}$ map $L$ into the right radical of $b$, which is zero, so they all vanish on $L$. Hence every $u\in\mathcal{S}$ induces a $\mathbb{D}$-linear mapping $\overline{u}:U/L\rightarrow V$, and we have $\overline{b}(\overline{x},\overline{u}(\overline{x}))=0$ for all $x\in U/L$. We denote by $\overline{\mathcal{S}}$ the space of all mappings $\overline{u}$ with $u\in\mathcal{S}$.
Conversely, for all $u\in\mathcal{A}_{\overline{b},\mathbb{D}}$ it is clear that we have $u\circ\pi\in\mathcal{S}$ for the standard projection $\pi:U\twoheadrightarrow U/L$. Therefore $\overline{\mathcal{S}}=\mathcal{A}_{\overline{b},\mathbb{D}}$. Since $u\in\mathcal{S}\mapsto\overline{u}\in\overline{\mathcal{S}}$ is clearly injective, we deduce that $\dim\mathcal{A}_{b,\mathbb{D}}=\dim\mathcal{A}_{\overline{b},\mathbb{D}}$.
@@
Denote by $P$ the Gram matrix of $B$ in the bases $U$ and $V$, and set $\mathcal{M}=\mathcal{S}\mathcal{H}_{n}(\mathbb{D})$. Then the elements of $\mathcal{S}$ are represented by the elements of $P^{-1}\mathcal{M}$ in the bases $\mathbf{B}_{U}$ and $\mathbf{B}_{V}$, and in particular $\dim\mathcal{A}_{b,\mathbb{D}}=\dim\mathcal{M}$. Then, it is easily checked in each case that
@@
Next, denote by $X$ the first vector of the standard basis of $\mathbb{D}^{n}$. Then we observe that $\mathcal{M}X=(\operatorname{\operatorname{Ker}}e)\times\mathbb{D}^{n-1}$ and hence $\dim_{\mathbb{F}}(\mathcal{S}x)=nd-1$. Moreover if $n\geq 2$ and we take $X^{\prime}$ as the second vector of the standard basis we find $\mathcal{M}X^{\prime}=\mathbb{D}\times(\operatorname{\operatorname{Ker}}e)\times\mathbb{D}^{n-2}$, and it follows that $\mathcal{S}x+\mathcal{S}x^{\prime}=V$ for some $x^{\prime}\in U$. If $d>1$ or $n>1$ we have collected enough information to see that $\mathcal{S}$ is target-reduced, and then we conclude from Lemma 5.3 that $\operatorname{\operatorname{Alt}}(\mathcal{S})=\mathbb{F}b$. If $d=1$ and $n=1$ then $\mathcal{S}=\{0\}$ and $\operatorname{\operatorname{Alt}}(\mathcal{S})$ has dimension $1$ (it is the space of all $\mathbb{F}$-bilinear forms on $U\times V$).
@@
  • Let $x\in V\smallsetminus\{0\}$. Then as $\dim_{\mathbb{F}}(\mathcal{S}x)=nd-1$ and $\mathcal{S}x$ is included in the right-orthogonal of $x$ under $b$, it equals this right-orthogonal, which yields $b(x,x)\neq 0$ since $x\not\in\mathcal{S}x$. Hence $b$ is nonisotropic.
  • Let us finally prove that $\mathcal{S}$ is irreducible. Let $x\in V\smallsetminus\{0\}$. Combining $\dim_{\mathbb{F}}(\mathcal{S}x)=nd-1$ with $x\not\in\mathcal{S}x$, we find $\mathbb{F}x\oplus\mathcal{S}x=V$. Therefore, $\mathcal{S}$ is irreducible.
6 Trivial spectrum subspaces
@@ 6.1 The Invariant Subspace Lemma
Let $\mathcal{S}$ be an $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$. Assume that there exists a $\mathbb{D}$-linear subspace $W$ of $V$ such that $\mathcal{S}$ contains every operator $u\in\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ that satisfies $\operatorname{\operatorname{Im}}u\subseteq W\subseteq\operatorname{\operatorname{Ker}}u$. Then $W$ is invariant under $\mathcal{S}$.
@@
Let $u\in\mathcal{S}$. Assume that there is a (nonzero) vector $x\in W$ such that $u(x)\not\in W$. The vector u(x)xu(x)-x is not in $W$, and hence it belongs to a direct summand $W^{\prime}$ of $W$ in $V$ (i.e., a $\mathbb{D}$-linear subspace of $V$ such that $V=W\oplus W^{\prime}$). Denote by $\pi$ the projection of $V$ onto $W$ along $W^{\prime}$, and by $v\in\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ the linear mapping that vanishes on $W$ and satisfies $v(y)=\pi(u(y))$ for all $y\in W^{\prime}$. Then $v\in\mathcal{S}$ by our assumptions, and hence $u-v\in\mathcal{S}$. Note that uvu-v leaves $W^{\prime}$ invariant. Moreover $(u-v)(x)=u(x)\equiv x$ mod $W^{\prime}$. The endomorphism of $V/W^{\prime}$ induced by uvu-v then has a nonzero fixed vector, which contradicts the assumption that it should have $\mathbb{F}$-trivial spectrum. Therefore $W$ is invariant under $\mathcal{S}$. ∎
@@ 6.2 The partial structure theorem for optimal spaces
Now, we shall complete the proof of Theorem 1.7. The case $n=1$ is obvious. So, we let $n\geq 2$ be an integer, $V$ be an $n$-dimensional $\mathbb{D}$-vector space, and $\mathcal{S}$ be an irreducible $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ with dimension $\alpha(n,d)$. We also assume that $|\mathbb{F}|\geq nd$.
We immediately note that $\mathcal{S}$ is intransitive.
Case 1. The space $\mathcal{S}$ is primitively intransitive.
Then $\mathcal{S}$ is deeply intransitive, by Proposition 4.5.
@@
  • Assume first that $(n,d)\neq(2,1)$. Then $n\geq\max(4-d,2)$, and hence $\dim\mathcal{S}\geq\alpha(n,d)-n+\max(4-d,2)$. By point (c) of Theorem 1.20, we find that $\mathcal{S}$ has a nondegenerate alternator $b$.
@@
  • Assume now that $(n,d)=(2,1)$. Then we simply note that $\dim\mathcal{S}=1$, so we can write $\mathcal{S}=\mathbb{F}u_{0}$ for some operator $u_{0}\in\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$; then $u_{0}$ is invertible because of the irreducibility assumption, and hence by picking an arbitrary symplectic form $s$ on the $\mathbb{F}$-vector space $V$ (which exists because $n=2$ and $d=1$), we see that $b:(x,y)\mapsto s(x,u_{0}^{-1}(y))$ is a nondegenerate alternator of $\mathcal{S}$.
@@
In any case, we have found a nondegenerate alternator $b$ of $\mathcal{S}$, to the effect that $\mathcal{S}\subseteq\mathcal{A}_{b,\mathbb{D}}$. Yet $\mathcal{A}_{b,\mathbb{D}}$ is deeply intransitive (by point (b) of Proposition 1.18), and hence $\dim\mathcal{A}_{b,\mathbb{D}}\leq\alpha(n,d)$ by point (a) of Theorem 1.20. Therefore $\mathcal{S}=\mathcal{A}_{b,\mathbb{D}}$.
Case 2. The space $\mathcal{S}$ is not primitively intransitive.
@@
Then we take a maximal $\mathbb{D}$-linear subspace $V^{\prime}$ of $V$ such that, for the standard projection $\pi:V\twoheadrightarrow V/V^{\prime}$, the projected space $\pi\mathcal{S}$ is intransitive. Note that $\{0\}\subset V^{\prime}\subset V$. By the maximality of $V^{\prime}$, the space $\pi\mathcal{S}$ is primitively intransitive, and hence deeply intransitive (see Proposition 4.5). Set then $t:=\dim_{\mathbb{D}}V^{\prime}$. Applying point (a) of Theorem 1.20, we find $\dim(\pi\mathcal{S})\leq\alpha(n-t,d)$. Now, we look at the space $\mathcal{S}^{V^{\prime}}$. First of all, the rank theorem yields
$\dim(\mathcal{S}^{V^{\prime}})=\dim(\mathcal{S})-\dim(\pi\mathcal{S})\geq\alpha(n,d)-\alpha(n-t,d)=dt(n-t)+\alpha(t,d).$
@@
Assigning to every $f\in\mathcal{S}^{V^{\prime}}$ its restriction $f_{V^{\prime}}\in\operatorname{\operatorname{End}}_{\mathbb{D}}(V^{\prime})$, we create an $\mathbb{F}$-linear mapping $\Phi$ whose range is an $\mathbb{F}$-trivial spectrum linear subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V^{\prime})$ and whose kernel is the subspace $\mathcal{D}$ of all $f\in\mathcal{S}^{V^{\prime}}$ that vanish on $V^{\prime}$. By Theorem 3.1 we have $\operatorname{\operatorname{rk}}\Phi\leq\alpha(t,d)$, whence the rank theorem yields
$\dim\mathcal{D}\geq\dim(\mathcal{S}^{V^{\prime}})-\alpha(t,d)\geq dt(n-t).$
@@
Yet $\mathcal{D}$ is included in the space of all $\mathbb{D}$-endomorphisms of $V$ that vanish on $V^{\prime}$ and map into $V^{\prime}$, a space which has dimension dt(nt)dt(n-t) over $\mathbb{F}$. Hence $\mathcal{D}$ equals that space, to the effect that all such endomorphisms belong to $\mathcal{S}$.
@@
Then the Invariant Subspace Lemma yields that $V^{\prime}$ is invariant under $\mathcal{S}$. However this contradicts the assumption that $\mathcal{S}$ is irreducible, and we conclude that Case 2 is actually impossible.

### v2

@@ 5.6 The converse statement
Let $U$ and $V$ be finite-dimensional $\mathbb{D}$-vector spaces, with $\dim_{\mathbb{D}}U=\dim_{\mathbb{D}}V>0$, and set $n:=\dim_{\mathbb{D}}U$. Let $S$ be an $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{Hom}}_{\mathbb{D}}(U,V)$ such that $\dim_{\mathbb{F}}(Sx)=nd-1$ for all $x\in U\smallsetminus\{0\}$. Assume that $S$ is target-reduced. Then all the nonzero alternators of $S$ are nondegenerate, and $\operatorname{\operatorname{Alt}}(S)$ has dimension at most $1$.
@@
Let $b\in\operatorname{\operatorname{Alt}}(S)$ be nonzero and degenerate. It follows that $nd=\dim_{\mathbb{F}}V>1$. Every element of $S$ maps the left radical $L$ of $b$ into its right radical $R$, and here $\dim_{\mathbb{F}}L=\dim_{\mathbb{F}}R$ since $\dim_{\mathbb{F}}U=\dim_{\mathbb{F}}V$.
We choose $x\in L\smallsetminus\{0\}$ and deduce that $\dim_{\mathbb{F}}R\geq\dim_{\mathbb{F}}Sx=nd-1$. Hence $\operatorname{\operatorname{rk}}(b)=1$. Then each bilinear form $(x,y)\in U^{2}\mapsto b(x,u(y))$, with $u\in S$, is alternating with rank at most $1$, and hence is zero. Hence $\operatorname{\operatorname{Im}}u\subseteq R$ for all $u\in S$, which contradicts the assumption that $S$ is target-reduced. Hence, every nonzero element of $\operatorname{\operatorname{Alt}}(S)$ is nondegenerate.
Next, let $b$ and $b^{\prime}$ belong to $\operatorname{\operatorname{Alt}}(S)\smallsetminus\{0\}$. Choose $x\in U\smallsetminus\{0\}$. By the first step, $b(x,-)\neq 0$. The nonzero linear forms $b^{\prime}(x,-)$ and b(x,)b(x,-) vanish on the $\mathbb{F}$-linear hyperplane $Sx$, and hence $b^{\prime}(x,-)=\lambda b(x,-)$ for some $\lambda\in\mathbb{F}$. Then $b^{\prime}-\lambda b$ is a degenerate element in $\operatorname{\operatorname{Alt}}(S)$, and by the first step $b^{\prime}-\lambda b=0$. We conclude that $\dim\operatorname{\operatorname{Alt}}(S)\leq 1$. ∎
@@
Set $S:=\mathcal{A}_{b,\mathbb{D}}$. To start with, to prove the first point we reduce the situation to the one where $B$ is nondegenerate. To see this, we consider the left radical $L$ of $B$, which equals the one of $b$, and the induced form $\overline{B}:(U/L)\times V\rightarrow\mathbb{D}$, and the $\mathbb{F}$-bilinear form $\overline{b}:(x,y)\in(U/L)\times V\mapsto e(\overline{B}(x,y))$. All the operators in $S$ map $L$ into the right radical of $b$, which is zero, so they all vanish on $L$. Hence every $u\in S$ induces a $\mathbb{D}$-linear mapping $\overline{u}:U/L\rightarrow V$, and we have $\overline{b}(\overline{x},\overline{u}(\overline{x}))=0$ for all $x\in U/L$. We denote by $\overline{S}$ the space of all mappings $\overline{u}$ with $u\in S$.
Conversely, for all $u\in\mathcal{A}_{\overline{b},\mathbb{D}}$ it is clear that we have $u\circ\pi\in S$ for the standard projection $\pi:U\twoheadrightarrow U/L$. Therefore $\overline{S}=\mathcal{A}_{\overline{b},\mathbb{D}}$. Since $u\in S\mapsto\overline{u}\in\overline{S}$ is clearly injective, we deduce that $\dim\mathcal{A}_{b,\mathbb{D}}=\dim\mathcal{A}_{\overline{b},\mathbb{D}}$.
@@
Denote by $P$ the Gram matrix of $B$ in the bases $U$ and $V$, and set $\mathcal{M}=S\mathcal{H}_{n}(\mathbb{D})$. Then the elements of $S$ are represented by the elements of $P^{-1}\mathcal{M}$ in the bases $\mathbf{B}_{U}$ and $\mathbf{B}_{V}$, and in particular $\dim\mathcal{A}_{b,\mathbb{D}}=\dim\mathcal{M}$. Then, it is easily checked in each case that
@@
Next, denote by $X$ the first vector of the standard basis of $\mathbb{D}^{n}$. Then we observe that $\mathcal{M}X=(\operatorname{\operatorname{Ker}}e)\times\mathbb{D}^{n-1}$ and hence $\dim_{\mathbb{F}}(Sx)=nd-1$. Moreover if $n\geq 2$ and we take $X^{\prime}$ as the second vector of the standard basis we find $\mathcal{M}X^{\prime}=\mathbb{D}\times(\operatorname{\operatorname{Ker}}e)\times\mathbb{D}^{n-2}$, and it follows that $Sx+Sx^{\prime}=V$ for some $x^{\prime}\in U$. If $d>1$ or $n>1$ we have collected enough information to see that $S$ is target-reduced, and then we conclude from Lemma 5.3 that $\operatorname{\operatorname{Alt}}(S)=\mathbb{F}b$. If $d=1$ and $n=1$ then $S=\{0\}$ and $\operatorname{\operatorname{Alt}}(S)$ has dimension $1$ (it is the space of all $\mathbb{F}$-bilinear forms on $U\times V$).
@@
  • Let $x\in V\smallsetminus\{0\}$. Then as $\dim_{\mathbb{F}}(Sx)=nd-1$ and $Sx$ is included in the right-orthogonal of $x$ under $b$, it equals this right-orthogonal, which yields $b(x,x)\neq 0$ since $x\not\in Sx$. Hence $b$ is nonisotropic.
  • Let us finally prove that $S$ is irreducible. Let $x\in V\smallsetminus\{0\}$. Combining $\dim_{\mathbb{F}}(Sx)=nd-1$ with $x\not\in Sx$, we find $\mathbb{F}x\oplus Sx=V$. Therefore, $S$ is irreducible.
6 Optimal trivial spectrum subspaces
@@ 6.1 The Invariant Subspace Lemma
Let $S$ be a trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$. Assume that there exists a $\mathbb{D}$-linear subspace $W$ of $V$ such that $S$ contains every operator $u\in\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ that satisfies $\operatorname{\operatorname{Im}}u\subseteq W\subseteq\operatorname{\operatorname{Ker}}u$. Then $W$ is invariant under $S$.
@@
Let $u\in S$. Assume that there is a (nonzero) vector $x\in W$ such that $u(x)\not\in W$. The vector u(x)xu(x)-x is not in $W$, and hence it belongs to a direct summand $W^{\prime}$ of $W$ in $V$ (i.e., a $\mathbb{D}$-linear subspace of $V$ such that $V=W\oplus W^{\prime}$). Denote by $\pi$ the projection of $V$ onto $W$ along $W^{\prime}$, and by $v\in\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ the linear mapping that vanishes on $W$ and satisfies $v(y)=\pi(u(y))$ for all $y\in W^{\prime}$. Then $v\in S$ by our assumptions, and hence $u-v\in S$. Note that uvu-v leaves $W^{\prime}$ invariant. Moreover $(u-v)(x)=u(x)\equiv x$ mod $W^{\prime}$. The endomorphism of $V/W^{\prime}$ induced by uvu-v then has a nonzero fixed vector, which contradicts the assumption that it should have trivial spectrum. Therefore $W$ is invariant under $S$. ∎
@@ 6.2 The partial structure theorem for optimal spaces
Now, we shall complete the proof of Theorem 1.9. The case $n=1$ is obvious. So, we let $n\geq 2$ be an integer, $V$ be an $n$-dimensional $\mathbb{D}$-vector space, and $S$ be an irreducible trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ with dimension $\alpha(n,d)$. We also assume that $|\mathbb{F}|\geq nd$.
We immediately note that $S$ is intransitive.
Case 1. The space $S$ is primitively intransitive.
Then $S$ is deeply intransitive, by Proposition 4.5.
@@
  • Assume first that $(n,d)\neq(2,1)$. Then $n\geq\max(4-d,2)$, and hence $\dim S\geq\alpha(n,d)-n+\max(4-d,2)$. By point (c) of Theorem 1.22, we find that $S$ has a nondegenerate alternator $b$.
@@
  • Assume now that $(n,d)=(2,1)$. Then we simply note that $\dim S=1$, so we can write $S=\mathbb{F}u_{0}$ for some operator $u_{0}\in\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$; then $u_{0}$ is invertible because of the irreducibility assumption, and hence by picking an arbitrary symplectic form $s$ on the $\mathbb{F}$-vector space $V$ (which exists because $n=2$ and $d=1$), we see that $b:(x,y)\mapsto s(x,u_{0}^{-1}(y))$ is a nondegenerate alternator of $S$.
@@
In any case, we have found a nondegenerate alternator $b$ of $S$, to the effect that $S\subseteq\mathcal{A}_{b,\mathbb{D}}$. Yet $\mathcal{A}_{b,\mathbb{D}}$ is deeply intransitive (by point (b) of Proposition 1.20), and hence $\dim\mathcal{A}_{b,\mathbb{D}}\leq\alpha(n,d)$ by point (a) of Theorem 1.22. Therefore $S=\mathcal{A}_{b,\mathbb{D}}$.
Case 2. The space $S$ is not primitively intransitive.
@@
Then we take a maximal $\mathbb{D}$-linear subspace $V^{\prime}$ of $V$ such that, for the standard projection $\pi:V\twoheadrightarrow V/V^{\prime}$, the projected space $\pi S$ is intransitive. Note that $\{0\}\subset V^{\prime}\subset V$. By the maximality of $V^{\prime}$, the space $\pi S$ is primitively intransitive, and hence deeply intransitive (see Proposition 4.5). Set then $t:=\dim_{\mathbb{D}}V^{\prime}$. Applying point (a) of Theorem 1.22, we find $\dim(\pi S)\leq\alpha(n-t,d)$. Now, we look at the space $S^{V^{\prime}}$. First of all, the rank theorem yields
$\dim(S^{V^{\prime}})=\dim(S)-\dim(\pi S)\geq\alpha(n,d)-\alpha(n-t,d)=dt(n-t)+\alpha(t,d).$
@@
Assigning to every $f\in S^{V^{\prime}}$ its restriction $f_{V^{\prime}}\in\operatorname{\operatorname{End}}_{\mathbb{D}}(V^{\prime})$, we create an $\mathbb{F}$-linear mapping $\Phi$ whose range is a trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V^{\prime})$ and whose kernel is the subspace $\mathcal{D}$ of all $f\in S^{V^{\prime}}$ that vanish on $V^{\prime}$. By Theorem 3.1 we have $\operatorname{\operatorname{rk}}\Phi\leq\alpha(t,d)$, whence the rank theorem yields
$\dim\mathcal{D}\geq\dim(S^{V^{\prime}})-\alpha(t,d)\geq dt(n-t).$
@@
Yet $\mathcal{D}$ is included in the space of all $\mathbb{D}$-endomorphisms of $V$ that vanish on $V^{\prime}$ and map into $V^{\prime}$, a space which has dimension dt(nt)dt(n-t) over $\mathbb{F}$. Hence $\mathcal{D}$ equals that space, to the effect that all such endomorphisms belong to $S$.
@@
Then the Invariant Subspace Lemma yields that $V^{\prime}$ is invariant under $S$. However this contradicts the assumption that $S$ is irreducible, and we conclude that Case 2 is actually impossible.